What is the confidence level of each of the following confidence intervals for the population mean μ?

i) ±1.96(σ/\sqrt{n})

ii) ±1.645(σ/\sqrt{n})

iii) ±2.575(σ/\sqrt{n})

iv) ± 1.28(σ/\sqrt{n})

v)  ±0.99(σ/\sqrt{n})

Thank You!
Answer #1

SOLUTION:

i) ±1.96(σ/\sqrt{n})

At 95% confidence level

\alpha = 1 - 95%  

\alpha = 1 - 0.95 =0.05

\alpha/2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

II)ii) ±1.645(σ/\sqrt{n})

At 90% confidence level

\alpha = 1 - 90%  

\alpha = 1 - 0.90 =0.10

\alpha/2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

3)

±2.575(σ/\sqrt{n})

At 99% confidence level the z is ,

\alpha  = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576

4)± 1.28(σ/\sqrt{n})

Z\alpha/2 = Z0.900 = 1.28

5)

±0.99(σ/\sqrt{n})

Z\alpha/2 = Z0.839 = 0.99